OpenStudy (anonymous):

Let f(x) = x2 - 8x + 5. Find f(-1).

OpenStudy (jadeishere):

plug in -1 for x (x-1)^2 - 8(-1) + 5 f(x)= -1 * -1 + 8 + 5 f(x) = 1 + 8 + 5 f(x) = 15

OpenStudy (jadeishere):

*14

OpenStudy (anonymous):

thank you

OpenStudy (jadeishere):

Yep :)